OpenStudy (anonymous):

Write the equation in standard form for the parabola. vertex (6,8) contains the point (4,-2) axis of symmetry x=6

OpenStudy (ageta):

little help @jhannybean

OpenStudy (sleepyhead314):

first get all the information given into an equation of a parabola doesn't matter which form for the first step can you do that? :) @rislam21

OpenStudy (sleepyhead314):

hint: vertex form for a parabola is y = a(x - h)^2 + k when (h , k) is the vertex

OpenStudy (anonymous):

Okay so I think that this parabola is going down, meaning it will be in the y-k=1/4p (x-h)² form and by using the coordinates of the vertex, the equation is now y-8=1/4p (x-h)² the problem is that I don't know how to get the value of p to finish the equation.

OpenStudy (anonymous):

(x-6)²**** sorry

OpenStudy (sleepyhead314):

hmmm I suggest plugging in the vertex first to get y = a (x - 6)^2 + 8 then to temporarily plug in the point (4, -2) for (x, y) to solve for a

OpenStudy (anonymous):

is this it: y= -14(x-6)² +8

OpenStudy (sleepyhead314):

hmmmm -2 = a (4 - 6)^2 + 8 can you solve for a from there? :/

OpenStudy (anonymous):

does a equal -2.5

OpenStudy (sleepyhead314):

yes! :D so plug a = -2.5 into y = a (x - 6)^2 + 8

OpenStudy (anonymous):

y= -2.5(x-6) ² +8 then y= x ² -12x +33.5

OpenStudy (sleepyhead314):

mmmm (x - 6)^2 = (x-6)(x-6) = x^2 - 12x + 36 -2.5(x^2 - 12x + 36) --> -2.5x^2 + 30x - 90 + 8 = ?

OpenStudy (anonymous):

ohh ok -2.5x ² +30x -82

OpenStudy (sleepyhead314):

yes! :D tahdahhh

OpenStudy (anonymous):

wait no that cant be the answer to my question because the standard form for a parabola where the vertex is not at the origin is in the form x-h=1/4p (y-k) ² if the parabola is going left or right and y-k=1/4p(x-h) ² if the parabola is going up or down

OpenStudy (sleepyhead314):

ohhhh ok whoops sorry >,< I haven't dont conics in a long time :/ axis of symmetry is x = 6 so going up or down so use y - k = 1/4 p(x - h)^2 plug in the vertex (6, 8) then do what we did before to try to solve for p?

OpenStudy (anonymous):

no problem! yea I don't know how to solve for P this is all I could do y-8=1/4p x^2 -12x +36 I don't know what to do next:(

OpenStudy (sleepyhead314):

good :) I think you should temporarily plug in the point (4 ,-2) for (x , y) to solve for p

OpenStudy (anonymous):

ahould I plug 4 into x^2 so that is become 16?

OpenStudy (sleepyhead314):

yes :)

OpenStudy (anonymous):

I got up to this: -14= 1/4p what do I do next? Should I divide both sides by 1/4?

OpenStudy (sleepyhead314):

multiply both sides by 4

OpenStudy (anonymous):

p= -56

OpenStudy (sleepyhead314):

actually... -2 - 8 = 1/4 p (16 - 48 + 36) -10 = 1/4 p (4) -10 = p

OpenStudy (anonymous):

Oh so the equation would be y-8= -1/40 (x-6)^2 ?

OpenStudy (sleepyhead314):

-10/4 = -5/2 y - 8 = -5/2 (x-6)^2

OpenStudy (anonymous):

wait how did you get -5/2

OpenStudy (sleepyhead314):

(1/4)p p = -10 (1/4) (-10) = -10/4

OpenStudy (anonymous):

Ohh okie dokie thanks so much for your help!!!!!!!!!!!!!

OpenStudy (sleepyhead314):

:)

OpenStudy (ageta):

cool